# MyFramelinecolor=black, outerlinewidth=2pt, roundcorner=20pt, innertopmargin=4pt, innerbottommargin=4pt, innerrightmargin=4pt, innerleftmargin=4pt, leftmargin = 4pt, rightmargin = 4pt, backgroundcolor=codeBackground

### v3

@@ 6.5.1 Properties of our Noise Matrices
Let $A$ be an $n\times n$ matrix (with $n<n$) whose entries are chosen independently and uniformly from the integer range $[-q,q]$. Then, $A$ is full rank with high probability, namely,
$P(\text{rank}(A)=n)=1-O\left(\frac{1}{q^{n-d+1}}\right).$
@@
The matrix $A$ will have rank $n$ if and only if its $n$ columns are linearly independent. For $A$ to be rank-deficient (rank $<d$), there must exist a non-trivial linear combination of the columns that results in the zero vector. This is equivalent to saying that one of the columns lies in the span of the preceding columns.
We can compute the probability that each new column is linearly independent of the previous ones as we proceed through the matrix construction column by column.
The first column of $A$ is non-zero with probability $1-\frac{1}{q^{n}}$, since it must not lie in the zero subspace. Given that the first column is non-zero, the second column is linearly independent of the first with probability $1-\frac{1}{q^{n-1}}$, as it must not lie in the one-dimensional subspace spanned by the first column. More generally, for the $k$-th column to be linearly independent of the previous $k-1$ columns, it must avoid the $(k-1)$-dimensional subspace spanned by those columns. This occurs with probability $1-\frac{1}{q^{n-k+1}}$.
The probability that all $n$ columns are linearly independent is the product of these probabilities, giving that
$P(\text{rank}(A)=n)=\prod_{k=0}^{n-1}\left(1-\frac{1}{q^{n-k}}\right)\geq\prod_{k=0}^{n-1}\exp\left({-\frac{1}{q^{n-k}}}\right)=\exp\left({-\sum_{k=0}^{d-1}\frac{1}{q^{n-k}}}\right),$
where the inequality holds since $\exp^{-x}<1-x$ for $x\in[0,1]$. Since the sum $\sum_{k=0}^{d-1}\frac{1}{q^{n-k}}$ is dominated by its last term, $\frac{1}{q^{n-d+1}}$, we get that
$\sum_{k=0}^{d-1}\frac{1}{q^{n-k}}=\Theta\left(\frac{1}{q^{n-d+1}}\right).$
Therefore, since $1-x+\frac{x^{2}}{2}<e^{-x}$ for $x\in[0,1]$, we conclude that
$P(\text{rank}(A)=d)\geq\exp\left({-\Theta\left(\frac{1}{q^{n-d+1}}\right)}\right)\geq 1-O\left(\frac{1}{q^{n-d+1}}\right).$
@@ 7 Using Native AI Workloads for Blockchain Consensus
The computational burden in AI models primarily arises from matrix multiplications. For matrices of size $n\times k$ and $k\times m$, the multiplication has a complexity of $nkm$. When $n=k=m$, this simplifies to $O(n^{3})$. Other operations, including activation functions, normalizations, rounding, etc. typically scale with $O(n^{2})$ as they operate element-wise.
@@ Acknowledgements
Appendix A A Self-Canceling Noise Scheme for high-Rank matrices
The reason we used additive noise (rather than multiplicative) in our PoUW Scheme (Algorithm 6.5) is in order to deal with the case where either $A$ or $B$ are the all-0 matrices (say), in which case any multiplicative operation is futile. Nevertheless, typical useful matrices are never of this form, so one observation is that we may w.l.o.g impose any statistical or algebraic property $\mathcal{P}(A)$ on $A$ (or $B$), so long as it is: (i) satisfied by real-world distributions with high probability $99.9\%$ (or else all bets are off and the protocol rejects); and (ii) efficiently provable and verifiable in sublinear time (ideally, one can implement a SNARK for it in practice). Assuming these conditions are met, restricting $A,B$ to the desired sub-family of matrices (i.e., passing the randomized property-test) opens up the possibility of using alternative—possibly more efficient—noise operators.
@@ Appendix A A Self-Canceling Noise Scheme for high-Rank matrices
In particular, we observe that using pseudorandom rotations on high-rank matrices, enables avoiding the decoding step in Algorithm 6.5 of “peeling" the noise altogether. The idea is to randomly rotate $A,B$ using a fast (pseudo-random) orthonormal matrix $R$, e.g., the FastJL Transform [AC09] which allows $O(n^{2}\log n)$-time rotations using the Fast Fourier Transform (FFT), and run the same PoUW scheme from Algorithm 6.5 on the product of
@@
The main point is that, while the output is preserved $A^{\prime}B^{\prime}=ARR^{\top}B=AB$ (no need to “‘peel-off” the noise), this operation completely “scrambles" the partial sums $P_{I,K,J}=\sum_{i\in I,k\in K,j\in J}(AR)_{ik}(R^{\top}B)_{kj}$ of intermediate $r\times r$ tiles ($|I|=|K|=|J|=r$). In particular, assuming the columns of $A$ and rows of $B$ have high rank (say full-rank $n$ for simplicity), the FastJL transofm $R$ guarantees that every (say) $\sqrt{n}\times\sqrt{n}$ tile of $AR$ is a random subspace embedding [CW13], so we may as well assume $A=B=I_{n}$. In this case, computing the partial sums $(RR^{\top})_{I,J,K}\in\mathbb{R}^{\sqrt{n}\times\sqrt{n}}$ seems hard, as it precludes exploiting the global (this time, FFT) structure of $R$ on submatrices – Indeed, it is not known how to multiply arbitrary $r\times r$ submatrices of the FFT matrix by general $x\in\mathbb{R}^{r}$ in $o(r^{2})$ time, unless massive pre-processing is allowed [KU11]. More formally, we conjecture that, for any full-rank $A,B\in\mathbb{R}^{n\times n}$, the amortized cost of computing all partial sums $(ARR^{\top}B)_{I,K,J}$ even conditioned on any (precomputed) function $f(A,B)$, still requires the trivial $\Theta\left(r^{3}\right)$ time (analogue of Assumption 6.4).
Efficient Proof of Proximity for Matrix Rank?
As mentioned earlier, the Pseudorandom-Rotation scheme requires an efficient SNARK (proof of proximity) for testing (or more modestly, distinguishing) high-from-low rank matrices $A,B$ ($\leq k$ vs. $\geq ck$) – In principle, this problem is equivalent to Low-Degree testing of the characteristic polynomial of $A$, which has a fast and practical SNARK ($O(\log^{2}k)$ [RS97, BSBHR18]), alas evaluating $p_{A}(x):=\det(A-x\cdot I)$ at arbitrary points $x$ explicitly has exponentially many terms, hence approximation via sampling seems necessary to make this idea work.
We leave the theoretical and practical exploration of the (pseudo-)random rotation scheme for future work.
@@ Appendix B Proof of Useful Work as a Poisson Process
Consider a sequence of $n$ independent and identically distributed Bernoulli random variables $E_{1},\ldots,E_{n}$ with parameter $p$. (That is, each $E_{i}$ is 1 with probability $p$ and 0 with probability 00.) The distribution of $\sum_{i=1}^{n}E_{i}$ follows a binomial distribution $B(n,p)$. It is known that the binomial distribution converges towards the Poisson distribution as $n\rightarrow\infty$ and $np$ converges to a finite limit (i.e., if $n$ is sufficiently large and $p$ is sufficiently small).
@@
{mdframed}
Honest mining process:
  1. 1.
    Initialize $i=1$.
  2. 2.
    Sample a uniformly random $\sigma_{i}$, and choose $\mathsf{x}_{i}$ and $\mathsf{data}_{i}$.
  3. 3.
    Compute $\pi_{i}=\mathsf{Solve}(\sigma_{i},\mathsf{data}_{i},\mathsf{x}_{i})$.
  4. 4.
    Compute $E_{i}=\mathsf{Verify}(\sigma_{i},\mathsf{data}_{i},\pi_{i})$.
  5. 5.
    Increase $i$ by 1 and return to step 2.
$(T,\overline{T},\alpha)$-adversarial mining process:
  1. 1.
    Run the honest process for $T$ steps resulting with $E_{1},\ldots,E_{T}$ and $\{(\sigma_{i},\mathsf{x}_{i},\mathsf{data}_{i},\pi_{i})\}_{i\in[T]}$.
  2. 2.
    Given all of the above, initialize (in polynomial time) a time $\alpha(n)\cdot(t(n)\cdot\overline{T})$ adversary. Upon a set of uniformly random $\sigma_{T+1},\ldots,\sigma_{T+\overline{T}}$, run it to generate $\pi_{T+1},\ldots,\pi_{T+\overline{T}}$. Compute $E_{i}=\mathsf{Verify}(\sigma_{i},\mathsf{data}_{i},\tilde{\pi}_{i})$ for $T<i\leq T+\overline{T}$.
  3. 3.
    Continue running the honest process with $i=T+\overline{T}+1$.
@@
We argue that a PoUW for blockchains, as in Definition B.3, along with standard properties of the cryptographic hash function, can be used to establish the security of the Bitcoin backbone protocol [GKL24]. More precisely, PoUW can be used to instantiate the same functionality as that of Bitcoin, i.e., a public ledger, as defined in [GKL24]. The proof follows the one presented in [GKP20], where we instantiate signatures of work with PoUW. We refer to [GKP20] for the proof.

### v4

@@ 6.5.1 Properties of our Noise Matrices
Let $A$ be an $n\times d$ matrix (with $d<n$) whose entries are chosen independently and uniformly from $\mathbb{Z}_{q}$. Then,
$P(\text{rank}(A)=d)=1-O\left({1/q^{n-d+1}}\right).$
@@
The matrix $A$ will have rank $d$ if and only if its $d$ columns are linearly independent. For $A$ to be rank-deficient (rank <d<d), there must exist a non-trivial linear combination of the columns that results in the zero vector. This is equivalent to saying that one of the columns lies in the span of the preceding columns.
The first column of $A$ is non-zero with probability $1-\frac{1}{q^{n}}$, since it must not lie in the zero subspace. Given that the first column is non-zero, the second column is linearly independent of the first with probability $1-\frac{1}{q^{n-1}}$, as it must not lie in the one-dimensional subspace spanned by the first column. More generally, for the $k$-th column to be linearly independent of the previous k1k-1 columns, it must avoid the (k1)(k-1)-dimensional subspace spanned by those columns. This occurs with probability $1-\frac{1}{q^{n-k+1}}$.
The probability that all $d$ columns are linearly independent is the product of these probabilities, giving that
$\Pr(\text{rank}(A)=n)=\prod_{k=0}^{d-1}\left(1-\frac{1}{q^{n-k}}\right)\geq\prod_{k=0}^{d-1}\exp\left({-\frac{1}{q^{n-k}}}\right)=\exp\left({-\sum_{k=0}^{d-1}\frac{1}{q^{n-k}}}\right),$
where the inequality holds since $e^{-x}<1-x$ for $x\in[0,1]$. Further, since $\sum_{k=0}^{d-1}\frac{1}{q^{n-k}}\leq 2/q^{n-d+1}$ and since $1-x+\frac{x^{2}}{2}<e^{-x}$ for $x\in[0,1]$, we conclude that $\Pr(\text{rank}(A)=n)\geq\exp\left({-{2}/{q^{n-d+1}}}\right)\geq 1-2/q^{n-d+1}.$
@@ 7 Using Native AI Workloads for Blockchain Consensus
The computational burden in AI models primarily arises from matrix multiplications. For matrices of size $n\times k$ and $k\times m$, the multiplication has a complexity of nkmnkm. When $n=k=m$, this simplifies to $O(n^{3})$. Other operations, including activation functions, normalizations, rounding, etc. typically scale with $O(n^{2})$ as they operate element-wise.
@@ Acknowledgements
Appendix A A Self-Canceling Noise Scheme
There is in principle nothing “holy” about the low-rank structure of the noise matrices used in the cuPOW  scheme (Algorithm 6.5) – Indeed, all we seem to need is a low-entropy perturbation operation $(A,B)\rightarrow(A^{\prime},B^{\prime})$, whose structure can be exploited for decoding the output $AB$, but not for intermediate computations – i.e., the perturbed matrices $A^{\prime},B^{\prime}$ should have marginally uniform projections on $r\times r$ tiles (which is the hard distribution for MatMul).
@@ Appendix A A Self-Canceling Noise Scheme
We observe that using pseudorandom rotations on high-rank matrices enables to avoid the decoding (“noise peeling") step in Algorithm 6.5 altogether. The idea is to randomly rotate $A,B$ using a fast pseudorandom orthonormal matrix $R$, e.g., the Randomized Hadamard Transform [AC06] which allows $O(n^{2}\log n)$-time rotations $AR$ using the Fast Fourier Transform (FFT), and run the same tile-based PoW scheme from Algorithm 6.5 cuPOW(A’,B’), on the product of
@@
The key point is that, while the matrices $RR^{\top}=I_{n}$ cancel at the output $A^{\prime}B^{\prime}=ARR^{\top}B=AB$ (hence no need to “‘peel-off” the noise), this cancellation does not occur locally on the partial sums
$P_{I,K,J}=\sum_{i\in I,k\in K,j\in J}(AR)_{ik}(R^{\top}B)_{kj}$ (A.1)
of intermediate $r\times r$ tiles ($|I|=|K|=|J|=r$).
Of course, this scheme, described in Algorithm A, is useless if the (dishonest) miner chooses the all-zero matrices $A=B=0^{n\times n}$. Nevertheless, assuming the columns of $A$ and rows of $B$ have high rank (which can be ensured by adding a random permutation of the identity matrix), say $\mathsf{rk}(A),\mathsf{rk}(B)\gtrsim n$ for simplicity), the FastJL transofm $R$ guarantees that every (say) $\sqrt{n}\times\sqrt{n}$ tile of $AR$ is a random subspace embedding [CW13], so we may as well assume $A=B=I_{n}$.
Local-vs.-Global Structure of FFT
Assuming $A=B=I_{n}$ for simplicity, the problem of computing the partial sums $(RR^{\top})_{I,J,K}\in\mathbb{R}^{r\times r}$ appears to be hard, since the global structure of the randomized Hadamard matrix seems hard to exploit locally on submatrices. Indeed, it is not known how to multiply arbitrary $r\times r$ submatrices of the FFT matrix by general $x\in\mathbb{R}^{r}$ in $o(r^{2})$ time, unless massive pre-processing is allowed [KU11].
To ensure the marginal entropy of every $r\times r$ tile of $(AR)\times(R^{\top}B)$ is full ($\sim r^{2}$ bits), Algorithm A composes the Hadamard transform with a block-diagonal matrix $D$ – choosing the block-size to be $d=r$ would ensure this, but this will lead to overall runtime of $O(n^{2}(r+\log n))$ of Algorithm A, no faster than the PoUW scheme from Algorithm 6.3 ($O(n^{2}r)$). Nevertheless, we conjecture that even block-size $d=O(1)$ remains computationally hard without massive preprocessing.
Conjecture A.1.
Let $A,B\in\mathbb{R}^{n\times n}$ be full-rank matrices. Then the amortized cost of computing all $n^{3}/r^{3}$ partial sums (A.1) in Algorithm A with $r=\sqrt{n},d=O(1)$, even with $O(n^{3})$ preprocessing time, requires $\Omega\left(r^{3}\right)$ time, in the word-RAM model with word-size $w=O(\log n)$.
This is the analogue of Assumption 6.4.
Random-Rotation PoUW Scheme (Self-Canceling Noise) $\mathsf{rrPoW}_{r,d}(\sigma,A,B)$: $\quad\sslash\quad A,B\in\mathbb{R}^{n\times n}$ 1. Parameters: $r\in\mathbb{N}$ (PoW tile-size);    $d\in\mathbb{N}$ (Hadamard block-size). 2. Parse $\sigma=D=\mathsf{Diag}(B_{1},B_{2},\dots,B_{n/d})$, where each block $B_{i}\in\{-1,1\}^{d\times d},\\ (B_{i})_{jk}\sim\text{Unif}(\{-1,1\})$. $\quad\\ \sslash D\in\{-1,1\}^{n\times n}$ is a random block-diagonal matrix with block-size $d$. 3. Let $R\leftarrow H_{n}\cdot D$, where $H_{n}\in\{\pm 1\}^{n}$ is the $n\times n$ Hadamard matrix [AC06]. 4. $A^{\prime}\leftarrow AR\;\;,\;\;B^{\prime}\leftarrow R^{\top}B$. 5. Run $\texttt{cuPOW}_{r}(A^{\prime},B^{\prime})$ (Algorithm 6.3). Algorithm 1.1
We leave additional theoretical and practical exploration of the (pseudo-)random rotation scheme for future work.
@@ Appendix B Proof of Useful Work as a Poisson Process
Consider a sequence of $n$ independent and identically distributed Bernoulli random variables $E_{1},\ldots,E_{n}$ with parameter $p$. (That is, each $E_{i}$ is 1 with probability $p$ and 0 with probability 0.) The distribution of $\sum_{i=1}^{n}E_{i}$ follows a binomial distribution $B(n,p)$. It is known that the binomial distribution converges towards the Poisson distribution as $n\rightarrow\infty$ and $np$ converges to a finite limit (i.e., if $n$ is sufficiently large and $p$ is sufficiently small).
@@
Honest mining process: 1. Initialize $i=1$. 2. Sample a uniformly random $\sigma_{i}$, and choose $\mathsf{x}_{i}$ and $\mathsf{data}_{i}$. 3. Compute $\pi_{i}=\mathsf{Solve}(\sigma_{i},\mathsf{data}_{i},\mathsf{x}_{i})$. 4. Compute $E_{i}=\mathsf{Verify}(\sigma_{i},\mathsf{data}_{i},\pi_{i})$. 5. Increase $i$ by 1 and return to step 2. $(T,\overline{T},\alpha)$-adversarial mining process: 1. Run the honest process for $T$ steps resulting with $E_{1},\ldots,E_{T}$ and $\{(\sigma_{i},\mathsf{x}_{i},\mathsf{data}_{i},\pi_{i})\}_{i\in[T]}$. 2. Given all of the above, initialize (in polynomial time) a time $\alpha(n)\cdot(t(n)\cdot\overline{T})$ adversary. Upon a set of uniformly random $\sigma_{T+1},\ldots,\sigma_{T+\overline{T}}$, run it to generate $\pi_{T+1},\ldots,\pi_{T+\overline{T}}$. Compute $E_{i}=\mathsf{Verify}(\sigma_{i},\mathsf{data}_{i},\tilde{\pi}_{i})$ for $T<i\leq T+\overline{T}$. 3. Continue running the honest process with $i=T+\overline{T}+1$.
@@
We argue that a PoUW for blockchains, as in Definition B.3, along with standard properties of the cryptographic hash function, can be used to establish the security of the Bitcoin backbone protocol [GarayKL24]. More precisely, PoUW can be used to instantiate the same functionality as that of Bitcoin, i.e., a public ledger, as defined in [GarayKL24]. The proof follows the one presented in [GarayKP20], where we instantiate signatures of work with PoUW. We refer to [GarayKP20] for the proof.
Appendix C Unpredictable Encodings
In Section 6.2 we introduced a notion termed transcript unpredictable encodings that suffices for a proof of useful work for matrix multiplication. Here we provide an even more generic abstraction that we believe could be of independent interest. We call this primitive an unpredictable encoding scheme (UES) which is a form of an “instance optimal” randomized encoding scheme with some new and incomparable properties to classical randomized encoding schemes. A UES is associated with a function $f\colon\{0,1\}^{n}\to\{0,1\}^{*}$ with a canonical algorithm Alg with worst-case running time $t(n)$, and it consists of two operations:
  • $\mathsf{Encode}(x;r)$: for an input $x$ this probabilistic procedure outputs an encoding $\tilde{x}$.
[…]
Correctness.
We require that $\mathsf{Eval}(\mathsf{Encode}(x))=f(x)$ for every $x$.
Running time.
The worst-case running time of $\mathsf{Encode}$ plus the running time of $\mathsf{Eval}$ is $t(n)\cdot(1+o(1))$.
Hardness.
It should be hard to compute the encoding of a given $x$ on a random $r$ in $\ll$ the running time of $\mathsf{Encode}$. That is, there is a constant $C$ such that for every $x$ and every time $t^{\prime}(n)$ algorithm $\mathcal{A}$, it holds that $\Pr_{r}[\mathcal{A}(x,r)=\mathsf{Encode}(x;r)]\leq C\cdot(t^{\prime}(n)/t(n))$.
Notice that the transcript unpredictability definition from Section 6.2 is a special case of the above where the encoding $\tilde{x}$ is essentially the transcript of the canonical matrix multiplication algorithm applied on the encoded input matrices.
Relation to randomized encodings.
The above definition resembles, at least syntactically, randomized encodings [DBLP:conf/focs/IshaiK00, DBLP:journals/siamcomp/ApplebaumIK06], and so we explain the differences. A randomized encoding allows to express an arbitrary computation, given by a function $f$ and input $x$, by a randomized representation $\hat{f}(x)$ whose distribution (viewed as a random variable) encodes $f(x)$, while revealing nothing else regarding $f$ and $x$. We do not require privacy of $x$ given $\tilde{x}$. On the other hand, we require a fine-grained hardness property and also an “optimal” efficiency property.